OpenStudy (anonymous):

What are the equations of the asymptotes for this hyperbola? x^2/25-y^2/4=1

OpenStudy (tkhunny):

a^2 = 25 ==> a = 5 b^2 = 4 ==> b = 2 That's about all you need.

OpenStudy (jhannybean):

\[\large \frac{x^2}{25}-\frac{y^2}{4}=1 \] so the slopes of the asymptotes of a hyperbola opening up and down are \[\large \pm \frac{b}{a}\]\(b=2 , a=5\)

OpenStudy (jhannybean):

opening left and right*

OpenStudy (anonymous):

Thank you!